what is the positive difference between the number of doughnuts in 50 dozen and number of slices of pizza (8 slices per pizza) in 80 pizzas?
Answer by rothauserc(4718)

total donuts = 50 *12 = 600
total pizza slices = 8 * 80 = 640
| 600 - 640 | = 40
positive difference is 40